OpenStudy (anonymous):

Prove: sec^4 theta - tan^4 theta = 2sec^2 theta -1

OpenStudy (paxpolaris):

\[LHS = \sec^4\theta-\tan^4\theta\]\[=\left( \sec^2\theta+\tan^2\theta \right)\left( \sec^2\theta-\tan^2\theta \right)\]

OpenStudy (paxpolaris):

can you take it from there......

OpenStudy (paxpolaris):

\[\sin^2\theta+\cos^2\theta = 1\] divide everything with \(\cos^2\theta\) gives you\[\implies \tan^2\theta +1=\sec^2\theta\ \dots\Huge\star\] using this equation you can see that \(\left( \sec^2\theta-\tan^2\theta \right)=1\) and \(\tan^2\theta\) can be replaced with \(\large\sec^2\theta-1\)

OpenStudy (anonymous):

so is it sec^2 -1

OpenStudy (anonymous):

(sec2theta-tan2theta)*(sec2theta+tan2theta) 1*(sec2theta+sec2theta-1) 2sec2theta-1